please help!

A 500 g object moving along the x-axis experiences a force given by the following equation: F=[10N - (2N/m)x]x, where x is measured in meters. The object’s velocity atx = 0 m is 24 m/s. What is the object’s velocity (in m/s) at x = 5 m?
Answer #1

first you find the equation for acceleration
mass m = 500g = 0.5 kg
F=[10N - (2N/m)x]x,
ma = [10N - (2N/m)x]x,
a = 20x - 4x^2

then you need to replace a as following.

uploaded image

then you need to substitute this with a in equation and solve the equation as following.

uploaded image

continue solving and you get the following result.

uploaded image

answered by: akinlo
Answer #2

mass m = 500g = 0.5 kg
F=[10N - (2N/m)x]x,
ma = [10N - (2N/m)x]x,
a = 20x - 4x2

and u = 24 m/s , x = 5 m , v= ?

from 3 rd equation of motion

v2 - u2 = 2ax

v2 = 242 +2x(20x-4x2 )

v2 = 576+(40x2 - 8x3 )

where x = 5 m

v2 = 576

v = 24 m/s

answered by: tory
Answer #3

F=[10N - (2N/m)x]x = ma

Here m = 500g = 0.5 Kg

Therefore

20x - 2x^2 = dv/dt

Threfore

v = 20xt - 2x^2t + 24

Put x = 5,we get

V(at x = 5) = 100t - 50t + 24

V(at x = 5) = 50t + 24

answered by: Lojyn
Answer #4

F=[10N - (2N/m)x]x = ma

Here m = 500g = 0.5 Kg

Therefore

20x - 2x^2 = dv/dt

Threfore

v = 20xt - 2x^2t + 24

Put x = 5,we get

V(at x = 5) = 100t - 50t + 24

answered by: ght